# High-Quality Axions from Higher-Form Symmetries in Extra Dimensions

### v1

@@ I Introduction
Such a description is naturally furnished in the language of generalized global symmetries [23]. In particular, the fate of the extra-dimensional axion’s low-energy shift symmetry may be fully understood from the higher-form symmetries of the extra-dimensional gauge theory (as well as their higher-group and non-invertible cousins)  [^1] A satisfying overview of the relevant generalized global symmetries is beyond the scope of this work. For excellent recent reviews germane to the current discussion, see e.g. [20, 49, 86, 87, 88].. The relevance of generalized symmetries to extra-dimensional axions is almost as old as the symmetries themselves [25, 26, 27, 28, 29, 20, 22]. Of particular pertinence to this paper are refs. [20, 22], which identified a connection between the extra-dimensional axion shift symmetry and a higher-form electric symmetry, and which pointed out that the extra-dimensional axion is a Nambu-Goldstone boson in the same sense that a photon is a Nambu-Goldstone boson of a higher-form symmetry. However, to the best of our knowledge, higher-form symmetries have yet to be used to comprehensively describe the extra-dimensional axion quality problem. In this paper, we enumerate the generalized global symmetries of the simplest extra-dimensional axion arising from the compactification of a 5D Abelian gauge theory, systematically classify the various effects that break these symmetries, and parameterize their impact on extra-dimensional solutions to the strong CP problem. The relative simplicity of this classification underlines the relevance of generalized global symmetries to physics beyond the Standard Model. (See refs. [30, 31, 32, 33, 34] for previous applications of generalized symmetries to particle phenomenology.)
We begin in Section II with a review of extra-dimensional axions from 5D gauge theories, the associated electric 1-form and magnetic 2-form symmetries, and their fates under reduction to 4D. In Section III we discuss the three principal ways of generating a potential for the extra-dimensional axion by breaking the electric 1-form symmetry that protects it, and we discuss how the axion is modified if the symmetry is gauged. Although the magnetic 2-form symmetry does not directly protect the axion potential, monopoles can induce effects that break the electric 1-form symmetry, as we describe in Section IV. In Section V we relate the breaking of the electric and magnetic higher-form symmetries to the anticipated size of effects arising in a theory of quantum gravity via various weak gravity conjectures. Our conclusions and future directions are summarized in Section VI.
@@ II.1 The Axion as a Kaluza-Klein Zero Mode
We adopt a minimal extra-dimensional axion model consisting of a one-form $U(1)$ gauge field $C$ and a one-form $SU(3)$ gauge field $A_{G}$ existing on an $\mathbb{R}^{1,3}\times S^{1}$ spacetime [18]. One can furthermore add a boundary to the extra-dimensional topology to avoid excess low energy degrees of freedom and for consistent coupling to the Standard Model (see e.g. refs. [19, 35, 36] where an orbifold is utilized). We emphasize that these more involved models share the same higher-form symmetry features pointed out in this work, but we keep to the simplest possible model as it is most instructive.
@@
where $e_{4}^{2}\equiv e_{5}^{2}/(2\pi R)$, where $\tilde{G}$ is the 4D two-form remnant of $G$, and where the dots represents couplings involving higher KK modes, which are irrelevant in the low energy limit of the theory. We have also left out the 4D gauge field component of $C$ and the scalar modes associated with $G$, both of which can be set to zero by a suitable choice of boundary conditions [19, 35].
@@ II.2 Associated Higher-Form Symmetries and their Dimensional Reduction
In 5D, the transformation associated with the electric one-form symmetry is a shift of the gauge field $C\rightarrow C+\Lambda$ where $\Lambda$ is a flat connection. The $\mathbb{R}^{1,3}$ piece of this shift clearly yields an identical 4D one-form symmetry for the 4D gauge field remnant of $C$, provided this remnant has not been set to zero. Crucially, the $S^{1}$ piece of the symmetry gives rise to the real number axion shift symmetry in 4D. This can be seen by simply plugging the shift into the definition of the axion (2) and carrying out the integral
@@
where the constant takes on values in the entire $[0,2\pi)$ interval. Alternatively, the connection between the one-form symmetry and axion shift symmetry can be seen by noticing that the electric current (4) reduces to the axion shift symmetry current $J_{s}=f^{2}{\star}{\rm d}\theta$ upon dimensional reduction. Hence, the axion real number zero-form shift symmetry has its higher-dimensional origin in the electric $U(1)_{e}^{(1)}$ one-form symmetry of $C$ [20]. This is a crucial observation: the 4D axion potential is protected by the axion real number shift symmetry. Since this shift symmetry is embedded in a one-form electric symmetry in extra-dimensional models, the quality of the axion can be understood in terms of the breaking and gauging of this electric one-form symmetry. For the axion to acquire a potential, the shift symmetry and hence the electric one-form symmetry of $C$ must be broken. But one-form symmetries are “harder to break” in the sense that 1) there are generally fewer ways to break them, and 2) explicit breaking has a lesser effect on IR physics compared to the breaking of ordinary zero-form symmetries. This latter point can be seen in several ways – e.g. by dimensional analysis, decompactifying an extra dimension, or by appeals to locality [38]. It can be more rigorously shown via action deformations in mean string field theory [39]. With these preliminary observations out of the way, we emphasize the central lesson of this work:
@@
The axion real number shift symmetry is embedded in an electric one-form symmetry. This symmetry is difficult to break, protecting the axion potential from dangerous corrections from UV physics.
@@
This statement should be contrasted with the usual appeal to gauge invariance. The higher-form symmetries perspective has a number of advantages. First, it is a more physical viewpoint since the electric one-form symmetry is a real symmetry of the theory, as opposed to gauge symmetry, which is a redundancy of description. Second, the one-form symmetry breaking viewpoint is on more equal footing with its PQ counterpart. For the PQ axion, the real number shift symmetry current descends directly from the current associated with the zero-form phase rotation symmetry of a complex scalar field $\phi$, given by $J_{p}=i{\star}(\phi{\rm d}\phi^{*}-\phi^{*}{\rm d}\phi)$ [5, 6, 7, 8]. Third, the higher-form symmetries viewpoint is much more powerful as an organizing principle. One can write down a host of additional gauge invariant operators in a theory of the form (1), but one finds that only a select few of these operators are capable of generating an axion potential. On the other hand, armed with the knowledge that it is the electric one-form symmetry that protects the axion potential, one only needs to write down the set of operators that break this symmetry. In fact, we find that whenever the symmetry is broken (and cannot be restored via a suitable field or current redefinition), the axion gets a potential, and vice versa. Using this organizing principle, we have naturally been led to all known ways to generate a potential for an extra-dimensional axion of the form (1), which we discuss in Section III. We have also been led to new avenues for potential generation, in particular via the interplay between the electric one-form symmetry, magnetic charges, and the magnetic two-form symmetry of $C$.
@@ III Modifying the Electric One-Form Symmetry
Magnetic Gauging ${\rm d}J_{e}=\frac{M}{2\pi}{\rm d}K$ $\mathbb{Z}_{M}^{(1)}$ $m_{\theta}=\frac{M}{2\pi}\,g_{4}\,e_{K,4}$
@@
Table 1: The four principal ways of modifying the electric $U(1)_{e}^{(1)}$ one-form symmetry of the 5D axion gauge field $C$. These are: (1) explicit breaking by electrically charged matter, discussed in Section III.1, (2) explicit breaking by gauging a magnetic two-form symmetry, discussed in Section III.2 (3) gauging the electric one-form symmetry, discussed in Section III.3, and (4) turning on an ABJ anomaly, discussed in Section III.4. Definitions of fields and constants are found in each respective section. For each type of one-form symmetry modification, the table shows how the current equation ${\rm d}J_{e}=0$ is modified, what (invertible) subgroup symmetry survives the modification, and what the resultant axion potential is. Additional potential contributions from monopoles, discussed in Section IV, fall under these four categories as well. Kinetic mixing terms and Euler-Heisenberg-type irrelevant operators do not break the one-form symmetry, as discussed in Section III.5.
@@ III.1 Electrically Charged Matter
This rough estimate shows the direct role that the instanton action plays in exponentially suppressing the effective axion potential. From the electric one-form symmetry point of view, where Wilson lines are the appropriate charge operators, it is the non-contractiblility of these wrapped worldlines that is responsible for breaking the one-form symmetry badly enough for there to be an effect on the axion potential in the IR. Their extended nature means they must come exponentially suppressed, and the fact that they can be completely self-enclosing when wrapped along the extra dimension is exactly the statement that the charge operators associated with a one-form symmetry in 5D become the charge operators of a zero-form symmetry in 4D. Note also the $2\pi/q$ periodicity of the potential, reflecting the surviving discrete $\mathbb{Z}_{q}^{(1)}$ one-form symmetry, which becomes a discrete $\mathbb{Z}_{q}^{(0)}$ zero-form shift symmetry for the axion in the IR.
@@ III.2 Gauging the Magnetic Two-Form Symmetry
By direct dimensional reduction, we find that eq. (16) becomes a mass equation for the axion, yielding $m_{\theta}=g_{4}e_{K,4}M/2\pi$ where $e_{K,4}^{2}\equiv e_{K}^{2}/2\pi R$. Another way to understand this mass it to directly dimensionally reduce (13), yielding the 4D axion action
@@
${\rm d}{\star}{\rm d}\theta=g_{4}^{2}e_{K,4}^{2}\frac{M^{2}}{4\pi^{2}}{\star}(\theta-\kappa)\,,$ (18)
where $\kappa$ is an arbitrary closed zero-form (which by definition is just an ordinary integration constant). This once again shows that the axion obtains a mass $m_{\theta}=g_{4}e_{K,4}M/2\pi$. Furthermore, the theory enjoys a surviving discrete $\mathbb{Z}^{(0)}_{M}$ zero-form symmetry, descending from the unbroken $\mathbb{Z}^{(1)}_{M}$ symmetry in the UV. To make this manifest, $\kappa$ can be taken to transform as $\kappa\rightarrow\kappa+2\pi\mathbb{Z}/M$ under the axion shift symmetry.
@@ III.4 Breaking by ABJ Anomaly
This current equation is the 5D equivalent of a 4D non-Abelian ABJ anomaly equation [62, 63]. Like for the magnetically gauged action discussed in Section III.2, one might attempt to define a refined conserved current $J_{e}^{\prime}=J_{e}-\frac{N}{8\pi^{2}}\mathrm{Tr}\left[A_{G}\wedge G\right]$, which is formally conserved, but which is not gauge invariant and hence not a genuine operator of the theory. However, the associated symmetry defect operator given by $U(\Sigma)=\exp\left[i\alpha\int_{\Sigma}J_{e}^{\prime}\right]$ is gauge invariant under both small and large gauge transformations as long as $\alpha=2\pi\mathbb{Z}/N$. This shows that the ABJ action (24) breaks the electric $U(1)^{(1)}_{e}$ one-form symmetry of $C$ down to a discrete $\mathbb{Z}_{N}^{(1)}$ one-form symmetry.
@@
One might wonder whether adding terms of the form (24), but with $G$ replaced with a $U(1)$ field strength could generate a potential for the axion – after all, such an action would break the electric $U(1)_{e}^{(1)}$ one-form symmetry. However, there are no $U(1)$ instantons on a topologically trivial 4D spacetime, hence there cannot be instantons in the IR of our theory [65, 66]. Nevertheless, if the theory contains magnetic monopoles, in line with what is mandated by the magnetic Weak Gravity Conjecture, closed monopole worldvolumes do indeed generate a potential via (24[67, 47, 68, 69]. We discuss this contribution in Section IV and other generalized Weak Gravity Conjecture constraints in Section V.
@@ IV Magnetic Charges
While the symmetry is broken, it is worth noting that a non-invertible symmetry remains, as studied in ref. [75]. We saw that a term of this form with a non-Abelian gauge field leads to an axion potential via gauge instantons in the IR. There are no $U(1)$ instantons on a topologically trivial 4D spacetime [65, 66], so one would naively be led to believe that the term (33) cannot generate a potential for the axion. However, it was found by ref. [69] that in 4D axion theories with a coupling of the form $\theta{\rm d}A\wedge{\rm d}A$, virtual monopoles magnetically charged under $A$ do induce an effective potential for the axion. This can also happen in 5D theories with a term of the form (33), which we now outline.

### v2

@@ I Introduction
Such a description is naturally furnished in the language of generalized global symmetries [23]. In particular, the fate of the extra-dimensional axion’s low-energy shift symmetry may be fully understood from the higher-form symmetries of the extra-dimensional gauge theory (as well as their higher-group and non-invertible cousins)  [^1] A satisfying overview of the relevant generalized global symmetries is beyond the scope of this work. For excellent recent reviews germane to the current discussion, see e.g. [20, 91, 50, 92, 93, 94].. The relevance of generalized symmetries to extra-dimensional axions is almost as old as the symmetries themselves [25, 26, 27, 28, 29, 20, 22]. Of particular pertinence to this paper are refs. [20, 22], which identified a connection between the extra-dimensional axion shift symmetry and a higher-form electric symmetry, and which pointed out that the extra-dimensional axion is a Nambu-Goldstone boson in the same sense that a photon is a Nambu-Goldstone boson of a higher-form symmetry. However, to the best of our knowledge, higher-form symmetries have yet to be used to comprehensively describe the extra-dimensional axion quality problem. In this paper, we enumerate the generalized global symmetries of the simplest extra-dimensional axion arising from the compactification of a 5D Abelian gauge theory, systematically classify the various effects that break these symmetries, and parameterize their impact on extra-dimensional solutions to the strong CP problem. The relative simplicity of this classification underlines the relevance of generalized global symmetries to physics beyond the Standard Model. (See refs. [30, 31, 32, 33, 34] for previous applications of generalized symmetries to particle phenomenology.)
We begin in Section II with a review of extra-dimensional axions from 5D gauge theories, the associated electric 1-form and magnetic 2-form symmetries, and their fates under reduction to 4D. In Section III we discuss the three principal ways of generating a potential for the extra-dimensional axion by breaking the electric 1-form symmetry that protects it, and we discuss how the axion is modified if the symmetry is gauged. Although the magnetic 2-form symmetry does not directly protect the axion potential, monopoles can induce effects that break the electric 1-form symmetry, as we describe in Section IV. In Section V we relate the breaking of the electric and magnetic higher-form symmetries to the anticipated size of effects arising in a theory of quantum gravity via various weak gravity conjectures and the Completeness Hypothesis. Our conclusions and future directions are summarized in Section VI.
@@ II.1 The Axion as a Kaluza-Klein Zero Mode
We adopt a minimal extra-dimensional axion model consisting of a one-form $U(1)$ gauge field $C$ and a one-form $SU(3)$ gauge field $A_{G}$ existing on an $\mathbb{R}^{1,3}\times S^{1}$ spacetime [18]. One can furthermore add a boundary to the extra-dimensional topology to avoid excess low energy degrees of freedom and for consistent coupling to the Standard Model (see e.g. refs. [19, 35, 36] where an orbifold is utilized). We emphasize that these more involved models share the same higher-form symmetry features pointed out in this work  [^2] For example, if the extra dimension is an $S^{1}/\mathbb{Z}_{2}$ orbifold, the survival of the axion zero mode forces us to pick even orbifold parity for $C_{5}$ so that $C_{5}(x,-y)=+C_{5}(x,y)$. This choice poses no obstacle to continuously dragging the symmetry defect operator associated with the electric one-form symmetry safely across the orbifold fixed points, keeping the electric one-form symmetry intact. This in turn rules out the possibility that localized zero mode mass terms can be placed on branes (see also ref. [95] where restrictions on such mass terms are discussed from a rather different perspective)., but we keep to the simplest possible model as it is most instructive.
@@
where $e_{4}^{2}\equiv e_{5}^{2}/(2\pi R)$, where $\tilde{G}$ is the 4D two-form remnant of $G$, and where the dots represent couplings involving higher KK modes, which are irrelevant in the low energy limit of the theory. We have also left out the 4D gauge field component of $C$ and the scalar modes associated with $G$, both of which can be set to zero by a suitable choice of boundary conditions [19, 35].
@@ II.2 Associated Higher-Form Symmetries and their Dimensional Reduction
In 5D, the transformation associated with the electric one-form symmetry is a shift of the gauge field $C\rightarrow C+\Lambda$ where $\Lambda$ is a flat connection. The $\mathbb{R}^{1,3}$ piece of this shift clearly yields an identical 4D one-form symmetry for the 4D gauge field remnant of $C$, provided this remnant has not been set to zero. Crucially, the $S^{1}$ piece of the symmetry gives rise to the continuous axion shift symmetry in 4D. This can be seen by simply plugging the shift into the definition of the axion (2) and carrying out the integral
@@
where the constant takes on values in the entire $[0,2\pi)$ interval. Alternatively, the connection between the one-form symmetry and axion shift symmetry can be seen by noticing that the electric current (4) reduces to the axion shift symmetry current $J_{s}=f^{2}{\star}{\rm d}\theta$ upon dimensional reduction. Hence, the continuous axion zero-form shift symmetry has its higher-dimensional origin in the electric $U(1)_{e}^{(1)}$ one-form symmetry of $C$ [20]. This is a crucial observation: the 4D axion potential is protected by the continuous axion shift symmetry. Since this shift symmetry is embedded in a one-form electric symmetry in extra-dimensional models, the quality of the axion can be understood in terms of the breaking and gauging of this electric one-form symmetry. For the axion to acquire a potential, the shift symmetry and hence the electric one-form symmetry of $C$ must be broken. But one-form symmetries are “harder to break” in the sense that 1) there are generally fewer ways to break them, and 2) explicit breaking has a lesser effect on infrared physics compared to the breaking of ordinary zero-form symmetries. This latter point can be seen in several ways – e.g. by dimensional analysis, decompactifying an extra dimension, or by appeals to locality [39]. It can be more rigorously shown via action deformations in mean string field theory [40]. With these preliminary observations out of the way, we emphasize the central lesson of this work:
@@
The continuous axion shift symmetry is embedded in an electric one-form symmetry. This symmetry is difficult to break, protecting the axion potential from dangerous corrections from UV physics.
@@
This statement should be contrasted with the usual appeal to gauge invariance. The higher-form symmetries perspective has a number of advantages. First, it is a more physical viewpoint since the electric one-form symmetry is a real symmetry of the theory, as opposed to gauge symmetry, which is a redundancy of description. Second, the one-form symmetry breaking viewpoint is on more equal footing with its PQ counterpart. For the PQ axion, the continuous shift symmetry current descends directly from the current associated with the zero-form phase rotation symmetry of a complex scalar field $\phi$, given by $J_{p}=i{\star}(\phi{\rm d}\phi^{*}-\phi^{*}{\rm d}\phi)$ [5, 6, 7, 8]. Third, the higher-form symmetries viewpoint is much more powerful as an organizing principle. One can write down a host of additional gauge invariant operators in a theory of the form (1), but one finds that only a select few of these operators are capable of generating an axion potential. On the other hand, armed with the knowledge that it is the electric one-form symmetry that protects the axion potential, one only needs to write down the set of operators that break this symmetry. In fact, we find that whenever the symmetry is broken (and cannot be restored via a suitable field or current redefinition), the axion gets a potential, and vice versa. Using this organizing principle, we have naturally been led to all known ways to generate a potential for an extra-dimensional axion of the form (1), which we discuss in Section III. We have also been led to new avenues for potential generation, in particular via the interplay between the electric one-form symmetry, magnetic charges, and the magnetic two-form symmetry of $C$.
@@ III Modifying the Electric One-Form Symmetry
Magnetic Gauging ${\rm d}J_{e}=\frac{M}{2\pi}{\rm d}K$ $\mathbb{Z}_{M}^{(1)}$ $m_{\theta}=e_{K,4}\,\frac{M}{2\pi f}$
@@
Table 1: The four principal ways of modifying the electric $U(1)_{e}^{(1)}$ one-form symmetry of the 5D axion gauge field $C$. These are: (1) explicit breaking by electrically charged matter, discussed in Section III.1, (2) explicit breaking by gauging a magnetic two-form symmetry, discussed in Section III.2 (3), gauging the electric one-form symmetry, discussed in Section III.3, and (4) turning on an ABJ anomaly, discussed in Section III.4. Definitions of fields and constants are found in each respective section. For each type of one-form symmetry modification, the table shows how the current equation ${\rm d}J_{e}=0$ is modified, what (invertible) subgroup symmetry survives the modification, and what the resultant axion potential is. Additional potential contributions from monopoles, discussed in Section IV, fall under these four categories as well. Kinetic mixing terms and Euler-Heisenberg-type irrelevant operators do not break the one-form symmetry, as discussed in Section III.5.
@@ III.1 Electrically Charged Matter
This rough estimate shows the direct role that the instanton action plays in exponentially suppressing the effective axion potential. From the electric one-form symmetry point of view, where Wilson lines are the appropriate charge operators, it is the non-contractibility of these wrapped worldlines that is responsible for breaking the one-form symmetry badly enough for there to be an effect on the axion potential in the infrared. Their extended nature means they must come exponentially suppressed, and the fact that they can be completely self-enclosing when wrapped along the extra dimension is exactly the statement that the charge operators associated with a one-form symmetry in 5D become the charge operators of a zero-form symmetry in 4D. Note also the $2\pi/q$ periodicity of the potential, reflecting the surviving discrete $\mathbb{Z}_{q}^{(1)}$ one-form symmetry, which becomes a discrete $\mathbb{Z}_{q}^{(0)}$ zero-form shift symmetry for the axion in the infrared.
@@ III.2 Gauging the Magnetic Two-Form Symmetry
By direct dimensional reduction, we find that eq. (16) becomes a mass equation for the axion, yielding $m_{\theta}=e_{K,4}M/2\pi f$ where $e_{K,4}^{2}\equiv e_{K}^{2}/2\pi R$. Another way to understand this mass is to directly dimensionally reduce (13), yielding the 4D axion action
@@
${\rm d}{\star}{\rm d}\theta=\frac{e_{K,4}^{2}}{f^{2}}\frac{M^{2}}{4\pi^{2}}{\star}(\theta-\kappa)\,,$ (18)
where $\kappa$ is an arbitrary closed zero-form (which by definition is just an ordinary integration constant). This once again shows that the axion obtains a mass $m_{\theta}=e_{K,4}M/2\pi f$. Furthermore, the theory enjoys a surviving discrete $\mathbb{Z}^{(0)}_{M}$ zero-form symmetry, descending from the unbroken $\mathbb{Z}^{(1)}_{M}$ symmetry in the UV. To make this manifest, $\kappa$ can be taken to transform as $\kappa\rightarrow\kappa+2\pi\mathbb{Z}/M$ under the axion shift symmetry.
@@ III.4 Breaking by ABJ Anomaly
This current equation is the 5D equivalent of a 4D non-Abelian ABJ anomaly equation [63, 64]. Like for the magnetically gauged action discussed in Section III.2, one might attempt to define a refined conserved current $J_{e}^{\prime}=J_{e}-\frac{N}{8\pi^{2}}\mathrm{Tr}\left[A_{G}\wedge G+\frac{2}{3}A_{G}\wedge A_{G}\wedge A_{G}\right]$, which is formally conserved, but which is not gauge invariant and hence not a genuine operator of the theory. However, the associated symmetry defect operator given by $U(\Sigma)=\exp\left[i\alpha\int_{\Sigma}J_{e}^{\prime}\right]$ is gauge invariant under both small and large gauge transformations as long as $\alpha=2\pi\mathbb{Z}/N$. This shows that the ABJ action (24) breaks the electric $U(1)^{(1)}_{e}$ one-form symmetry of $C$ down to a discrete $\mathbb{Z}_{N}^{(1)}$ one-form symmetry.
@@
One might wonder whether adding terms of the form (24), but with $G$ replaced with a $U(1)$ field strength could generate a potential for the axion – after all, such an action would break the electric $U(1)_{e}^{(1)}$ one-form symmetry. However, there are no $U(1)$ instantons on a topologically trivial 4D spacetime, hence there cannot be instantons in the infrared regime of our theory [66, 67]. Nevertheless, if the theory contains magnetic monopoles, in line with what is mandated by the magnetic Weak Gravity Conjecture, closed monopole worldvolumes do indeed generate a potential via (24[68, 48, 69, 70]. We discuss this contribution in Section IV and other generalized Weak Gravity Conjecture constraints in Section V.
@@ IV Magnetic Charges
While the symmetry is broken, it is worth noting that a non-invertible symmetry remains, as studied in ref. [76]. We saw that a term of this form with a non-Abelian gauge field leads to an axion potential via gauge instantons in the infrared. There are no $U(1)$ instantons on a topologically trivial 4D spacetime [66, 67], so one would naively be led to believe that the term (33) cannot generate a potential for the axion. However, it was found by ref. [70] that in 4D axion theories with a coupling of the form $\theta{\rm d}A\wedge{\rm d}A$, virtual monopoles magnetically charged under $A$ do induce an effective potential for the axion  [^7] Ref. [29] linked this effective potential in 4D to the further breakdown of the axion shift symmetry, which would have a surviving non-invertible symmetry remnant were monopoles not included.. This can also happen in 5D theories with a term of the form (33), which we now outline.